OpenStudy (anonymous):

If the quadratic formula is used to find the solution set of 3x^2 + 4x - 2 = 0, what is the result of substituting in the values? I got -4 (plus or minus) the sqrt of 8 over 6

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[b^2-4ac=4^2-4\times 3\times (-2)\]

OpenStudy (anonymous):

that is \(16+24=40\)

OpenStudy (anonymous):

so -4(plus or minus) the sqrt of 40 over 6?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but you are not actually done

OpenStudy (anonymous):

that's all it has in my options

OpenStudy (anonymous):

it didn't say to go any further

OpenStudy (anonymous):

ok then you ARE done